OpenStudy (anonymous):

Evaluate 7+77+777+7777........upto n terms

OpenStudy (anonymous):

I am stuck on a step

OpenStudy (anonymous):

I got this

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Shall i tell my approach i am stuck on a step

ganeshie8 (ganeshie8):

yea please

OpenStudy (anonymous):

\[\frac{ 7 }{ 9 }\left\{ 9 +99 +999+....... \right\}\] \[\frac{ 7 }{ 9 }\left\{ (10-1) + (10^{2}-1)+..... \right\}\]

ganeshie8 (ganeshie8):

that looks great ! keep going

OpenStudy (anonymous):

No i could have but i am stuck , i mean how should i procedd

ganeshie8 (ganeshie8):

group the exponents and 1's separately

ganeshie8 (ganeshie8):

and keep in mind that there are n terms

ganeshie8 (ganeshie8):

\[\frac{ 7 }{ 9 }\left\{ (10-1) + (10^{2}-1)+..... \right\}\] \[\frac{ 7 }{ 9 }\left\{ (10 + 10^2 + \cdots 10^n) - (1 + 1 + \cdots \text{n terms})\right\}\]

ganeshie8 (ganeshie8):

\[\frac{ 7 }{ 9 }\left\{ (10 + 10^2 + \cdots 10^n) - n\right\}\]

OpenStudy (anonymous):

if they are similar we can write them as "one" Like n here?

ganeshie8 (ganeshie8):

1 + 1 = 2 1 + 1 + 1 = 3 1 + 1 + 1 + 1 = 4 1 + 1 + 1 + 1 + ... n times = n

OpenStudy (anonymous):

now that can be done thanks a ton

OpenStudy (anonymous):

can u plz help me after @ganeshie8